40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 28 03:21:35 2010, Suresh, Update, Locking, Aligned the MC2 transmission photodiode 
    Reply  Thu Oct 28 04:23:38 2010, rana, Update, IOO, A2L prep 
Message ID: 3804     Entry time: Thu Oct 28 03:21:35 2010     Reply to this: 3806
Author: Suresh 
Type: Update 
Category: Locking 
Subject: Aligned the MC2 transmission photodiode 

Yuta and Suresh

The MC2 transmission is seen on the QPD

Once the laser was locked to the cavity, and the PMC was able to follow the laser (ref: elogs by Yuta and Rana today)  we had a steady TEMoo mode in the MC.  This gave us sufficient transmission through MC2 to be easily visible with an IR viewer and we were able to guide the beam on to the QPD.  The beam however seemed to over fill the QPD, a lens (f=180mm) was placed before the beam folding mirror and the spot sized reduced.   The the QPD was found to be not fixed to the table and this was also recitified.  The signal level is still low: total signal is just about 7 DAQ steps amounting to about 5mV.  Tomorrow we plan to work on the alignment of the PSL and MC and thus increase this signal.

A new channel to observe the length variations in the MC.

A long BNC cable was laid from the MC locking electronics next (southwards) to the PSL table to the DAQ cards picking up the signals from the PRM OSEMS.  This is to pick up one of the MC locking signals and collect it on a DAQ channel.  However as there are no spare DAQ channels currently available one of the PRM OSEM (UL and LL) photodiode channels was unplugged and replaced with the signal from the long BNC cable.   For identifying the correct DAQ channel we put in a 2 Vpp 10Hz signal with a function generator into this BNC.  Tow signals can be picked up in this fashion and they are available on PRM_LLSEN_IN1 and PRM_ULSEN_IN1. We plan to use this for improving the alignment of the MC tomorrow. 

The algorithm for this alignment is that if the beam from the PSL is not centered on the MC1 then tilting MC1 would result in a change in the length of the cavity as seen by the light.  Using this as feedback the spot could be precisely centered on the MC1 and then the MC alignment could be completed by moving MC2 and MC3 to reobtain TEM_oo within the cavity.

ELOG V3.1.3-